OpenStudy (taepoohwae):

solve the equation and check the solution -8(t-8)-(t+8)=8 the solution is t=

OpenStudy (helpsos):

t = 48/9

OpenStudy (helpsos):

one sec ill write the solution for you....

OpenStudy (taepoohwae):

ok

OpenStudy (helpsos):

-8 (t -8)- (t + 8) = 8 -8t + 64-t - 8 = 8 -9t + 56 = 8 -9t = 8-56 -9t = -48 t = -48/-9 which can also be 48 /9 :)

OpenStudy (helpsos):

does that make sense?

OpenStudy (taepoohwae):

Lost me with the 9

OpenStudy (helpsos):

you mean the -48/ -9 ?

OpenStudy (helpsos):

or the -9t ?

OpenStudy (taepoohwae):

9t

OpenStudy (helpsos):

oh ok, there i combined like terms. There was -8t and a -t. when you put those two together you get -9t

OpenStudy (taepoohwae):

ok thanks

OpenStudy (helpsos):

np :) is there anything else?

OpenStudy (taepoohwae):

not at the moment. your help was greatly appreciated!

OpenStudy (helpsos):

thanx :)